# TreeKV: Smooth Key-Value Cache Compression with Tree Structures

### v1

@@ 5.2 Language Modeling
FollowingXiao et al. (2023); Han et al. (2024); Zhang et al. (2024c), we also thoroughly examine TreeKV to test if a LLM pre-trained with a 4k context window can effectively perform language modeling task on exceptionally extended text. We concatenate all the 100 books in PG19 test set to create a 10M length example and evaluate its generation capability using Llama-2-7B with StreamingLlm, H2O, TOVA, and TreeKV. Figure 4 shows the NLL curves for input lengths from 0.1M to 10M. The curves reveal that TOVA and H2O suffer performance degradation with longer sequences, leading to significantly worse negative log-likelihoods than StreamingLlm and TreeKV. In contrast, TreeKV consistently outperforms all other baseline methods including StreamingLlm, demonstrating superior capability with longer inputs. This phenomenon also clarifies why our method lags behind TOVA by 0.02 at a length of 4k but regains an advantage at lengths of 8k and beyond.
@@ 5.3 Long Context Understanding
Table 3 summarizes the performance metrics for all tasks and cache sizes in the Longbench benchmark. Overall, our model achieved an average improvement of 0.65 over H2O and 0.42 over SnapKV, demonstrating its superior capacity to retain and recall relevant information from extended text, it did not surpass the full attention model, which leaves space for us to improve on with minimal resource requirements.
@@ 6 Conclusion
In this paper, we first explore the frequency representations of information collected during generation using multi-level wavelet decomposition, leading to the development of TreeKV, a training-free method that utilizes a tree structure for smooth cache compression. Our evaluation shows that TreeKV outperforms existing methods in both prefilling and generation stages.

### v2

@@ 5.2 Language Modeling
Following Xiao et al. (2023); Han et al. (2024); Zhang et al. (2024c), we also thoroughly examine TreeKV to test if a LLM pre-trained with a 4k context window can effectively perform language modeling task on exceptionally extended text. We concatenate all the 100 books in PG19 test set to create a 10M length example and evaluate its generation capability using Llama-2-7B with StreamingLLM, H2O, TOVA, and TreeKV. Figure 4 shows the NLL curves for input lengths from 0.1M to 10M. The curves reveal that TOVA and H2O suffer performance degradation with longer sequences, leading to significantly worse NLL than StreamingLLM and TreeKV. In contrast, TreeKV consistently outperforms all other baseline methods including StreamingLLM, demonstrating superior capability with longer inputs.
@@ 5.3 Long Context Understanding
Table 3 summarizes the performance metrics for all tasks and cache sizes in the Longbench benchmark. Overall, our model achieved an average improvement of 0.74 over H2O and 0.48 over SnapKV, demonstrating its superior capacity to retain and recall relevant information from extended text, it did not surpass the full attention model, which leaves space for us to improve on with minimal resource requirements.
@@ 6 Conclusion
In this paper, we first explore the frequency representations of information collected during generation using multi-level wavelet decomposition, leading to the development of TreeKV, a training-free method that utilizes a tree structure for smooth cache compression. Our evaluation shows that TreeKV outperforms existing methods in both prefilling and generation setups.